OpenStudy (anonymous):

Find the area of the region that lies inside both curves r = A*sin(theta) and r = B*sin(theta), A > 0, B > 0?

OpenStudy (nikvist):

\[r_1=A\sin\theta\quad\Rightarrow\quad x^2+\left(y-\frac{A}{2}\right)^2=\left(\frac{A}{2}\right)^2\]\[r_2=B\sin\theta\quad\Rightarrow\quad x^2+\left(y-\frac{B}{2}\right)^2=\left(\frac{B}{2}\right)^2\]\[S=\frac{\pi}{4}\cdot\left(\min{\left\{A,B\right\}}\right)^2\]

OpenStudy (perl):

this does not work

OpenStudy (perl):

try A = 2 , B = 1

OpenStudy (perl):

should be 3pi/2 , not sure how you got your answer

OpenStudy (perl):

ok looks like you did a polar conversion r = A sin theta r^2 = r* A sin theta x^2 + y^2 = A r* sin theta x^2 + y^2 = Ay x^2 + y^2 - Ay = 0 complete square x^2 + (y -A/2)^2 = (A/2)^2 similiarly x^2 + (y-B/2)^2 = (B/2)^2 directly integrating polar is much easier, btw so now you have two cartesian equations , and you did ... ?

OpenStudy (perl):

im not sure how to to integrate it using cartesian equations, hmmm

OpenStudy (perl):

ok i guess it can be done, but it is tricky

OpenStudy (perl):

wlog A > B > 0 so

OpenStudy (nikvist):

These are circles. Why do you integrate?

OpenStudy (perl):